OpenStudy (anonymous):

Gabrielle is making a quilt. She purchases a piece of material 1 1/3 yd by 4 2/3 yd. (a) Gabrielle wants to use all the fabric to make quilt squares. What is the side length in inches of the largest square piece she can cut so that no material is left over? Show your work. Please: NO ANSWERS - please show the idea of how it is solved ONLY

OpenStudy (anonymous):

give me a min my internet is slow

OpenStudy (anonymous):

Ok thats fine

OpenStudy (perl):

I would first convert these to improper fractions

OpenStudy (perl):

1 1/3 = 1 + 1/3 = 4/3 yards = 1.333 yards 4 2/3 = 4 + 2/3 = 14/3 yards = 4.6667 yards

OpenStudy (anonymous):

I GOT IT!!!

OpenStudy (anonymous):

Ugh nevermind

OpenStudy (anonymous):

@Mendicant_Bias

OpenStudy (anonymous):

@jenheaven

OpenStudy (anonymous):

ok wht do u need help with

OpenStudy (anonymous):

same problem cuz i dont fully understand it yet

OpenStudy (anonymous):

well wht dont u understand

OpenStudy (anonymous):

I thought the answer would be 4 2/3 divided by 1 1/3 but I got 3 1/2. I dont know what to do with that number now

OpenStudy (anonymous):

wht bout simplify

OpenStudy (anonymous):

ing the problem

OpenStudy (anonymous):

3 and a half IS the answer but i dont know how its part of the square (if its the number of squares, the size of the square, or a randomnumber)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

???

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

i think it's the number of squares

OpenStudy (anonymous):

im not tht good at this type of math

OpenStudy (anonymous):

me neither :( I hate fractions.

OpenStudy (anonymous):

@MyaDaBoo

OpenStudy (anonymous):

Sorry i don't know how to do this :/

OpenStudy (anonymous):

:(